OpenStudy (anonymous):

Help!!!!!! Iris has been studying an invasive population of snails. This particular snail has no local predators so the population grows wildly. She has observed that the population follows an exponential rate of growth for fifteen years. 1. Create your own exponential function, f(x), which models the snail population. Explain to Iris how your function shows the principal population and the rate of growth, in complete sentences.

OpenStudy (anonymous):

@owl

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@ayoubEpst

OpenStudy (anonymous):

@lucyv7129

OpenStudy (anonymous):

@Official_Hotboy

OpenStudy (anonymous):

Here is some ideas to think about for this problem. First consider that the formula f(x) = P (1 + r) ^x. is the formula used for compound interest. in a bank if you invest $1 and the APR (annual percent rate) is 5% will you will end up with $1,05 at the end of the year ? ... YES, NO, MAYBE ??? The first consideration in this case is that you have to get at least your dollar back ... That is where the "1 +" comes in to play... now for you extra money (the interest) that is the "r" In my stort it would be 5% (.05)... so if you keep th emoney in there for one yeat you get 1(1+.05)^1 or 1.05 ... now the compounding ... if you keep it there for a second yeat then the total is 1(1.05)(1.05) OR .... 1(1.05)^2 Now if the compounding is more than once a year (lets say quarterly) then the compounding periods has to be multiplied by 4 and the rate has to be divided by 4 1(1 + .05/4)^ 2*4 for two years... the number of compoundings can be any number... Now ther is a SPECIAL situation where there is a infinite compoundings per year (and NO you do not end up with an infinite amount of money) when done the equation produces a special number "e" (2.71...) this is called continuoue compounding... this actually will model things like radioactive decay f(x)= P e^r bottom line here is if the snals reproduce continuously you would use the "e" formula... f(x) = P (1 + r) ^x. where r is the percent increase in population each cycle and x is the time interval of each cycle if the reproduction is at a steady interval then you can use the

OpenStudy (anonymous):

What is thus? @david111

OpenStudy (anonymous):

This*

OpenStudy (anonymous):

What do you need help with...? Choose some numbers and put them in the exponential growth formula.

OpenStudy (anonymous):

What is the exponential growth formula

OpenStudy (anonymous):

@david111

OpenStudy (anonymous):

\[F(x)=n _{o}e ^{rt}\]

OpenStudy (anonymous):

n0 = initial population r = rate of growth

OpenStudy (anonymous):

What does e stand for? Can you give me an example

OpenStudy (anonymous):

e is the number e.

OpenStudy (anonymous):

50% = 50/100 =0.5

OpenStudy (anonymous):

That would be e?

OpenStudy (anonymous):

0.5 means 50% growth per year. Initial is 5 snails.

OpenStudy (anonymous):

Can I do 5^15t

OpenStudy (anonymous):

F(x) = 15(.3)^x

OpenStudy (anonymous):

Can you explain this to me

OpenStudy (anonymous):

f(x) = P (1 + r) ^x.

OpenStudy (anonymous):

What is p?

OpenStudy (anonymous):

lets get strait that f(x) or f of anything just represents the y. your function is basically y=mx+b m- number of population increase in terms of people x- number of years b- number of people that are now present y- the total number of people depending on m,x and b.

OpenStudy (anonymous):

Okay so can I do fx= 100x+200

OpenStudy (anonymous):

@ganeshie8 @phi @amistre64 @themaster @mathmale @hoblos @satellite73 @hartnn @loser66 @agent0smith

OpenStudy (anonymous):

@david111 I'm begging you I have to finish this tonight please

OpenStudy (anonymous):

@ganeshie8 so how do I solve thus

OpenStudy (anonymous):

put f(x)= instead of N=

OpenStudy (anonymous):

That's the entire question

OpenStudy (anonymous):

I can't install the tea weaver I am on a ipad

OpenStudy (anonymous):

It's the first one